OpenStudy (anonymous):

√(8x+1)=5 show the steps please :)

OpenStudy (saifoo.khan):

\[\sqrt{8x+1}=5\]The square on both sides. \[8x+1 = 25\] Solve. =)

OpenStudy (unklerhaukus):

\[8x+1=25\]\[8x=24\]\[x=3\]